# Modeling the solar system as an Observed Multi-Transit System I: Characterization Limits from Analytic Timing Variations

### v3

@@ Abstract
software: CalcEph(Gastineau et al., 2015), TTVFaster (Agol & Deck, 2016), matplotlib (Hunter, 2007; Caswell et al., 2019), Julia (Bezanson et al., 2017), DataFrames.jl(Bouchet-Valat & Kamiński, 2023)
1 Introduction
@@
An opportunity for accurate characterization of an exoplanet system exists in a system that contains at least two transiting exoplanets. In a multi-transiting planet system (MTS), gravitationally-induced planet-planet interactions can significantly perturb the transit times (Ragozzine & Holman, 2018). Agol et al. (2005) and Holman & Murray (2005) independently recognized that the resulting transit timing variations (TTVs) could be inverted to measure masses of perturbed, near-resonant terrestrials (see also Miralda‐Escude, 2002; Schneider, 2004). Moreover, the TTV analysis can reveal non-transiting planets that were previously undetected (Ballard et al., 2011; Nesvorný et al., 2012), and may allow us to probe relatively large orbital distances.
@@
In the following section, we describe our procedure for simulating (§ 2.1) and fitting our suite of synthetic observations (§ 2.2). In the first half of section 3, we demonstrate our methodology for a fiducial case. We finish with a discussion of our overall results, and the impact of our findings.
2 Methods
@@
We then model transit times using TTVFaster (Agol & Deck, 2016, hereafter AD16), and afterwards determine how accurately the masses and orbital parameters of the planets can be recovered – including those whose properties can only be indirectly measured (i.e. Mars and Jupiter). This model assumes the planets are co-planar and edge-on in their orbits, which is appropriate as the RMS inclination of the eight solar system planets is less than 2(Winn & Fabrycky, 2015).
For each simulation, we perform a likelihood profile analysis to search for non-transiting objects, and globally fit the TTVs using the Levenberg-Marquardt algorithm (LM; Levenberg, 1944; Marquardt, 1963). Next, we retrieve the masses and orbital elements of the planets by carrying out an affine-invariant Markov Chain Monte Carlo (MCMC) analysis (Goodman & Weare, 2010). Using both the likelihood profile and the MCMC sampling has benefits: the former checks for the modality of the probability distribution, and the latter characterizes the model parameter uncertainties. In the following sections, we describe our methods in more detail.
2.1 Procedure for simulating data
2.1.1 Simulation specifications
The JPL ephemerides have been generated by fitting numerically integrated orbits to both ground-based and space-based observations of SS objects. For this work, we adopt the DE440 version of the ephemerides to compute planetary position, velocity, and acceleration vectors at a given Julian Date (JD; Park et al., 2021). Since a randomly-oriented, distant observer is unable to see all of the planets in the system transiting the Sun (Brakensiek & Ragozzine, 2016), we assume that the observer lies along a line defined by the intersections of Venus and the EMB’s orbital planes. Over numerous orbits, the location from which both Venus and EM could be seen transiting the Sun – dubbed the transit zone – would change very slightly due to planet-moon interactions, and planet-planet interactions (Heller & Pudritz, 2016). However, the individual transit zones for the terrestrials would be stable for thousands of years (Wells et al., 2017), and the timescales of our simulations are much shorter than secular timescales.
@@ 2.1.2 Mid-transit times
2.1.3 Injected timing noise
@@
2.2 Fitting transit times
@@ 2.2.1 Transit ephemeris
where $t_{0,i}$ is the first transit time, and $E$ is the number of orbits elapsed after the initial transit time (i.e. epoch). For each planet, we fit this linear model to the noisy transit times, and derive an estimate for the mean ephemeris ($t_{0,i}$, $P_{i}$). Table 2 shows expected transit times of the linear ephemeris, and deviations from the mean period model for Venus and EM. In practice, transit-timing data often have missing or discontinuous observations, while in this paper we simulate contiguous transits based on the computed orbits. This observation was simulated for a 30-yr timescale with Gaussian uncertainties at the 30-s level. This dataset will be our fiducial case, which we will closely examine in sections 3.13.3.
@@ 2.2.3 Maximum likelihood
2.2.4 Test configurations and likelihood profile
@@
  • We adopt the value corresponding to $\ln\mathcal{L}_{\text{max}}$ from the grid search (§ 2.2.4).
@@
2.3 Error analysis
Mathematically, Bayes’ theorem states that $p(\textbf{x}|\textbf{y})=p(\textbf{y}|\textbf{x})p(\textbf{x})/p(\textbf{y})$ (Bayes & Price, 1763). In terms of the likelihood, we can rewrite this equation to give the posterior probability distribution, $p(\textbf{x},\sigma_{\text{sys}})\propto\Pi(\textbf{x})\times\mathcal{L}(\textbf{x},\sigma_{\text{sys}})$ where $\Pi(\textbf{x})\equiv p(\textbf{x})$ is the prior function. While our approach to the likelihood in Eq. 8 does not account for the prior probability distribution, it provides an initial estimate with which to sample the posterior probability.
@@ 2.3 Error analysis
Using an affine-invariant Markov chain algorithm (Goodman & Weare, 2010), we sample our fitted parameter set and the systematic uncertainty, which is assumed to be the same for both planets. This parameter absorbs any inaccuracies in our model resulting in a posterior probability, $p(\textbf{x},\sigma_{\text{sys}})$. Given that high dimensional Markov chains require a large number of steps, we use $\approx 10^{4}$ steps with 75 random walkers. After the burn-in – which we define as the first time the walkers cross the median parameter (Knutson, 2009) – the chain is examined to assess convergence and calculate the minimum effective sample size [^4] Using https://github.com/TuringLang/MCMCDiagnosticTools.jl/.
3 Results
In the following subsections, we compare how well the test configurations describe the TTVs. Then, we explore the distributions of the retrieved masses and dynamics (§ 3.3), and the effect of timing precision and observing baseline (§ 3.4).
@@ 3 Results
3.1 Example simulations and fits
@@
3.2 The search for additional planets
@@
We summarize our sampling results for the fiducial case in Table 5. For each model, we report the median posterior parameters and a 68.3% quantile interval from the Markov chain analysis. We compute the Bayesian Inference Criterion (BIC; Schwarz, 1978; Kass & Raftery, 1995) with the $\chi^{2}$ that corresponds to the maximum value of Eqn. 8. The BIC estimates the Bayesian evidence for a given model by assuming that the posterior obeys the Laplace approximation. This assumption is not valid for all the posteriors (see the eccentricity vectors in Figure 7), but we can still use BIC to penalize models with additional free parameters. With respect to the maximum likelihood,
@@ 3.2.2 Three planets
Figure 15 displays a corner plot of the mass-ratios, periods, and eccentricity-vectors of this configuration. For the inner planets, the masses and periods follow Gaussian distributions, while the constraints on the eccentricity vectors are strongly correlated. We attribute this aspect to the eccentricity-eccentricity degeneracy described in Lithwick et al. (2012). We also find that planet d (i.e. Jupiter) has a negatively skewed e-vector; and the $e_{d}\cos{\omega_{d}}$-$\mu_{d}$ panel demonstrates a banana-shaped probability distribution function. This approximate degeneracy between masses and orbit shapes is known to impact measurement of planet masses due to observations with insufficient timing precision (see Hadden & Lithwick, 2017; Leleu et al., 2023).
3.2.3 Four planets
@@
Figure 7: 1D posterior histograms for planet parameters fit in at least two of our configurations for our fiducial case – i.e. those common between our 2-, 3-, and 4-planet models discussed in § 3.2 – compared to the truth (dashed vertical line) for Venus, Earth + Moon, and Jupiter. The rows correspond to planets b/c/d; the parameter being plotted is indicated in the title of each panel.
@@
3.3 Masses and orbital elements
@@
3.4 Effect of timing noise and observing baseline
@@
In contrast to the transiting exoplanets, the gas giant’s parameters appear much less accurate and precise in the 3-planet model. Upon visual examination of the traces – and employing the Gelman, Rubin, and Brooks diagnostic (Brooks & Gelman, 1998) – we found that the Markov chains did not actually converge for survey baselines shorter than 22 years (for 60 seconds of injected noise) and 23 years (when $\sigma_{obs}=90$). Therefore, we could not directly compare the percent errors for all the datasets.
@@
4 Discussion
4.1 Model Accuracy
@@
4.2 Implications for observational capabilities
@@
Realistically, one could measure the RV signals caused by a long-period giant planet which does not transit, if RV data are sufficiently precise (Pepe & Lovis, 2008). This would complement and confirm a transit-timing analysis, and vice versa. Moreover, a joint fit with simulated RV monitoring of the Sun could improve mass constraints for the outer gas giant, and may reduce the observing baseline required to accurately measure the masses of Venus and EM.
@@
4.3 Future work
In its current state, TTVFaster breaks down in accuracy at high eccentricities, for large planet-star mass ratios, and for planets close to resonance (Agol & Deck, 2016). Therefore, future work could involve implementing a more accurate model, such as the NBodyGradient code (see Agol et al., 2021) to analyze the simulated solar system. Similarly, a full photo-dynamical simulation of realistic transits with varying levels of white noise and correlated noise appropriate for our Sun could help us determine how precisely the times can be measured (a la Morris et al., 2020, but with bigger glass or multiple wavelengths to achieve higher precision). We leave investigating the impact of missed transits and transit-timing outliers to future work.
5 Conclusion
@@
Our focus was to compute the simultaneous detectability of an additional non-transiting gas giant and terrestrial planet, given detectable TTV signals from two transiting rocky planets. Our computed transit times are fully available online, along with the code used to analyze them [^6] https://github.com/bmlindor/ttv_ss. The results are as follows:
@@ Acknowledgements
This material is based upon work supported by the NSF Graduate Research Fellowship Program under Grant No. DGE-1762114. E.A. acknowledges support from NSF grant AST-1907342, NASA NExSS grant No. 80NSSC18K0829, and NASA XRP grant 80NSSC21K1111. This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program.
We thank Steven G. Johnson for creating packages to access Python functions and modules from the Julia language. We thank Matthew Hollman, Stephen Kane, and Brett Morris for useful discussions. We also thank the reviewers for their thoughtful comments which helped to improve this manuscript.
@@ Appendix A Posteriors
We provide the MCMC parameter median and $1\sigma$ confidence intervals from fitting the transit times of Venus/Earth across the Sun over 30 years. This dataset had $\sigma_{\mathrm{obs}}=30$ second Gaussian noise injected (discussed in §3.2). Figures 15 and 16 show corner plots of parameters in the 3-planet, and 4-planet models reported in Table 5. In the table’s right-most column, we provide masses and Keplerian orbital elements for Venus, Earth + Moon, Jupiter and Mars – adopted from Hussmann et al. (2009) or calculated using JPL ephemerides (Park et al., 2021).

### v4

# Modeling the Solar System I: Characterization Limits from Analytic Transit Timing Variations

@@ Abstract
software: CalcEph(Gastineau et al., 2015), TTVFaster (Agol and Deck, 2016), matplotlib (Hunter, 2007; Caswell et al., 2019), Julia (Bezanson et al., 2017), DataFrames.jl(Bouchet-Valat and Kamiński, 2023)
I Introduction
@@
An opportunity for accurate characterization of an exoplanet system exists in a system that contains at least two transiting exoplanets. In a multi-transiting planet system (MTS), gravitationally-induced planet-planet interactions can significantly perturb the transit times (Ragozzine and Holman, 2018). Agol et al. (2005) and Holman and Murray (2005) independently recognized that the resulting transit timing variations (TTVs) could be inverted to measure masses of perturbed, near-resonant terrestrials (see also Miralda‐Escude, 2002; Schneider, 2004). Moreover, the TTV analysis can reveal non-transiting planets that were previously undetected (Ballard et al., 2011; Nesvorný et al., 2012), and may allow us to probe relatively large orbital distances.
@@
In the following section, we describe our procedure for simulating (§ II.1) and fitting our suite of synthetic observations (§ II.2). In the first half of section III, we demonstrate our methodology for a fiducial case. We finish with a discussion of our overall results, and the impact of our findings.
II Methods
@@
We then model transit times using TTVFaster (Agol and Deck, 2016, hereafter AD16), and afterwards determine how accurately the masses and orbital parameters of the planets can be recovered – including those whose properties can only be indirectly measured (i.e. Mars and Jupiter). This model assumes the planets are co-planar and edge-on in their orbits, which is appropriate as the RMS inclination of the eight solar system planets is less than 2(Winn and Fabrycky, 2015).
For each simulation, we perform a likelihood profile analysis to search for non-transiting objects, and globally fit the TTVs using the Levenberg-Marquardt algorithm (LM; Levenberg, 1944; Marquardt, 1963). Next, we retrieve the masses and orbital elements of the planets by carrying out an affine-invariant Markov Chain Monte Carlo (MCMC) analysis (Goodman and Weare, 2010). Using both the likelihood profile and the MCMC sampling has benefits: the former checks for the modality of the probability distribution, and the latter characterizes the model parameter uncertainties. In the following sections, we describe our methods in more detail.
II.1 Procedure for simulating data
II.1.1 Simulation specifications
The JPL ephemerides have been generated by fitting numerically integrated orbits to both ground-based and space-based observations of SS objects. For this work, we adopt the DE440 version of the ephemerides to compute planetary position, velocity, and acceleration vectors at a given Julian Date (JD; Park et al., 2021). Since a randomly-oriented, distant observer is unable to see all of the planets in the system transiting the Sun (Brakensiek and Ragozzine, 2016), we assume that the observer lies along a line defined by the intersections of Venus and the EMB’s orbital planes. Over numerous orbits, the location from which both Venus and EM could be seen transiting the Sun – dubbed the transit zone – would change very slightly due to planet-moon interactions, and planet-planet interactions (Heller and Pudritz, 2016). However, the individual transit zones for the terrestrials would be stable for thousands of years (Wells et al., 2017), and the timescales of our simulations are much shorter than secular timescales.
@@ II.1.2 Mid-transit times
II.1.3 Injected timing noise
@@
II.2 Fitting transit times
@@ II.2.1 Transit ephemeris
where $t_{0,i}$ is the first transit time, and $E$ is the number of orbits elapsed after the initial transit time (i.e. epoch). For each planet, we fit this linear model to the noisy transit times, and derive an estimate for the mean ephemeris ($t_{0,i}$, $P_{i}$). Table 2 shows expected transit times of the linear ephemeris, and deviations from the mean period model for Venus and EM. In practice, transit-timing data often have missing or discontinuous observations, while in this paper we simulate contiguous transits based on the computed orbits. This observation was simulated for a 30-yr timescale with Gaussian uncertainties at the 30-s level. This dataset will be our fiducial case, which we will closely examine in sections III.1III.3.
@@ II.2.3 Maximum likelihood
II.2.4 Test configurations and likelihood profile
@@
  • We adopt the value corresponding to $\ln\mathcal{L}_{\text{max}}$ from the grid search (§ II.2.4).
@@
II.3 Error analysis
Mathematically, Bayes’ theorem states that $p(\textbf{x}|\textbf{y})=p(\textbf{y}|\textbf{x})p(\textbf{x})/p(\textbf{y})$ (Bayes and Price, 1763). In terms of the likelihood, we can rewrite this equation to give the posterior probability distribution, $p(\textbf{x},\sigma_{\text{sys}})\propto\Pi(\textbf{x})\times\mathcal{L}(\textbf{x},\sigma_{\text{sys}})$ where $\Pi(\textbf{x})\equiv p(\textbf{x})$ is the prior function. While our approach to the likelihood in Eq. 8 does not account for the prior probability distribution, it provides an initial estimate with which to sample the posterior probability.
@@ II.3 Error analysis
Using an affine-invariant Markov chain algorithm (Goodman and Weare, 2010), we sample our fitted parameter set and the systematic uncertainty, which is assumed to be the same for both planets. This parameter absorbs any inaccuracies in our model resulting in a posterior probability, $p(\textbf{x},\sigma_{\text{sys}})$. Given that high dimensional Markov chains require a large number of steps, we use $\approx 10^{4}$ steps with 75 random walkers. After the burn-in – which we define as the first time the walkers cross the median parameter (A., 2009) – the chain is examined to assess convergence and calculate the minimum effective sample size [^4] Using https://github.com/TuringLang/MCMCDiagnosticTools.jl/.
III Results
In the following subsections, we compare how well the test configurations describe the TTVs. Then, we explore the distributions of the retrieved masses and dynamics (§ III.3), and the effect of timing precision and observing baseline (§ III.4).
@@ III Results
III.1 Example simulations and fits
@@
III.2 The search for additional planets
@@
We summarize our sampling results for the fiducial case in Table 5. For each model, we report the median posterior parameters and a 68.3% quantile interval from the Markov chain analysis. We compute the Bayesian Inference Criterion (BIC; Schwarz, 1978; Kass and Raftery, 1995) with the $\chi^{2}$ that corresponds to the maximum value of Eqn. 8. The BIC estimates the Bayesian evidence for a given model by assuming that the posterior obeys the Laplace approximation. This assumption is not valid for all the posteriors (see the eccentricity vectors in Figure 7), but we can still use BIC to penalize models with additional free parameters. With respect to the maximum likelihood,
@@ III.2.2 Three planets
Figure 15 displays a corner plot of the mass-ratios, periods, and eccentricity-vectors of this configuration. For the inner planets, the masses and periods follow Gaussian distributions, while the constraints on the eccentricity vectors are strongly correlated. We attribute this aspect to the eccentricity-eccentricity degeneracy described in Lithwick et al. (2012). We also find that planet d (i.e. Jupiter) has a negatively skewed e-vector; and the $e_{d}\cos{\omega_{d}}$-$\mu_{d}$ panel demonstrates a banana-shaped probability distribution function. This approximate degeneracy between masses and orbit shapes is known to impact measurement of planet masses due to observations with insufficient timing precision (see Hadden and Lithwick, 2017; Leleu et al., 2023).
III.2.3 Four planets
@@
Figure 7: 1D posterior histograms for planet parameters fit in at least two of our configurations for our fiducial case – i.e. those common between our 2-, 3-, and 4-planet models discussed in § III.2 – compared to the truth (dashed vertical line) for Venus, Earth + Moon, and Jupiter. The rows correspond to planets b/c/d; the parameter being plotted is indicated in the title of each panel.
@@
III.3 Masses and orbital elements
@@
III.4 Effect of timing noise and observing baseline
@@
In contrast to the transiting exoplanets, the gas giant’s parameters appear much less accurate and precise in the 3-planet model. Upon visual examination of the traces – and employing the Gelman, Rubin, and Brooks diagnostic (Brooks and Gelman, 1998) – we found that the Markov chains did not actually converge for survey baselines shorter than 22 years (for 60 seconds of injected noise) and 23 years (when $\sigma_{obs}=90$). Therefore, we could not directly compare the percent errors for all the datasets.
@@
IV Discussion
IV.1 Model Accuracy
@@
IV.2 Implications for observational capabilities
@@
Realistically, one could measure the RV signals caused by a long-period giant planet which does not transit, if RV data are sufficiently precise (Pepe and Lovis, 2008). This would complement and confirm a transit-timing analysis, and vice versa. Moreover, a joint fit with simulated RV monitoring of the Sun could improve mass constraints for the outer gas giant, and may reduce the observing baseline required to accurately measure the masses of Venus and EM.
@@
IV.3 Future work
In its current state, TTVFaster breaks down in accuracy at high eccentricities, for large planet-star mass ratios, and for planets close to resonance (Agol and Deck, 2016). Therefore, future work could involve implementing a more accurate model, such as the NBodyGradient code (see Agol et al., 2021) to analyze the simulated solar system. Similarly, a full photo-dynamical simulation of realistic transits with varying levels of white noise and correlated noise appropriate for our Sun could help us determine how precisely the times can be measured (a la Morris et al., 2020, but with bigger glass or multiple wavelengths to achieve higher precision). We leave investigating the impact of missed transits and transit-timing outliers to future work.
V Conclusion
@@
Our focus was to compute the simultaneous detectability of an additional non-transiting gas giant and terrestrial planet, given detectable TTV signals from two transiting rocky planets. Our computed transit times are fully available online, along with the code used to analyze them [^6] https://doi.org/10.5281/zenodo.17156392 (catalog DOI: 10.5281/zenodo.17156392); https://github.com/bmlindor/ttv_ss. The results are as follows:
@@ Acknowledgements
This material is based upon work supported by the NSF Graduate Research Fellowship Program under Grant No. DGE-1762114. E.A. acknowledges support from NSF grant AST-1907342, NASA NExSS grant No. 80NSSC18K0829, and NASA XRP grant 80NSSC21K1111. This work was performed in part by the Virtual Planetary Laboratory Team, which is a member of the NASA Nexus for Exoplanet System Science, and funded via the NASA Astrobiology Program ICAR Grant 80NSSC23K1398. This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program.
We thank Steven G. Johnson for creating packages to access Python functions and modules from the Julia language. We thank Matthew Holman, Stephen Kane, and Brett Morris for useful discussions. We also thank the reviewers for their thoughtful comments which helped to improve this manuscript.
@@ Appendix A Posteriors
We provide the MCMC parameter median and $1\sigma$ confidence intervals from fitting the transit times of Venus/Earth across the Sun over 30 years. This dataset had $\sigma_{\mathrm{obs}}=30$ second Gaussian noise injected (discussed in §III.2). Figures 15 and 16 show corner plots of parameters in the 3-planet, and 4-planet models reported in Table 5. In the table’s right-most column, we provide masses and Keplerian orbital elements for Venus, Earth + Moon, Jupiter and Mars – adopted from Hussmann et al. (2009) or calculated using JPL ephemerides (Park et al., 2021).